# Multi-Virtual Knot Theory

### v1

@@ Abstract
We generalize virtual knot theory to multi-virtual knot theory where there are a multiplicity of virtual crossings. Each virtual crossing type can detour over the other virtual crossing types, and over the classical or flat classical crossings. New invariants of multi-virtual knots and links are introduced and new problems that arise are described. We show how the extensions of the Penrose coloring evaluation for trivalent plane graphs and our generalizations of this to non-planar graphs and arbitrary numbers of colors acts as a motivation for the construction of the multi-virtual theory.
@@ 1 Introduction
generalizations. Here we indicate a circled virtual crossing in the Penrose expansion. Sometimes the self-intersection is written without the circle, but the next paragraph explains why we have to indicate it. One can prove that is the number of proper three-colorings of a planar graph
Our approach to the Penrose bracket and its generalizations is based on the concept of a formation due to G. Spencer-Brown [53]. See Figure 17 for an illustration of a formation. In the initial case of three colors, we choose to illustrate directly the colors red (r) and blue (b) and to indicate the third color purple (p) as a superposition of red and blue. If one makes a planar drawing using red and blue Jordan curves so that the red curves are disjoint from one another, and the blue curves are disjoint from one another, but red and blue curves can share a finite set of disjoint intervals, then the resulting drawing can be interpreted as a trivalent graph with a proper three coloring. The edges colored purple are in correspondence with the intervals in the drawing shared by red and. blue. An examination of the figure should clarify this statement. In the figure the corresponding trivalent graph is shown and the edges of the graph that correspond to shared red and blue are labeled with a slash. These purple edges in the coloring correspond to a perfect matching in the trivalent graph. This illustrates how a proper coloring gives rise to an even perfect matching (See Section 2). Conversely, if one has a properly colored trivalent graph, then there is a formation that corresponds to the coloring. The red and blue curves are regarded as crossing one another when one curve starts on one side of the other, shares an interval, and continues on the other side of the other curve.
@@
The dotted crossing is regarded as a coloring vertex that demands that all four lines incident to this vertex have the same color. When a boxed or circled crossing has colors assigned to its arcs it can be regarded as two straight arcs that cross one another at the circled or boxed intersection. Each straight arc has one color. The two colors may be the same or different. These interpretations are important for understanding the relationships with colorings, but are not used in the formal expansions of the formulas themselves, except at the very end. The straight arcs of the circled or boxed flat crossing can be colored with two colors that are different or the same. The technical reasons why this method adjusts the Penrose bracket count are reviewed in Section 2. Here it is important to point out that this second recursion is defined on the states of the Penrose bracket expansion. Each state consists in a collection of planar Jordan curves intersecting in round and square virtual crossings. The above evaluation is equivalent to a graphical contraction deletion algorithm for a graph where each state loop corresponds to a node of and two nodes are connected by an edge only when their loops touch at a boxed virtual crossing. Round virtual crossings do not affect the graph Then the above relation becomes the contraction deletion relation shown below.
@@
This means that we evaluate a dichromatic polynomial [37] for each state, and that the generalized Penrose evaluation is a sum with coefficients over these dichromatic evaluations. This is a useful direction of reformulation and will be followed up in subsequent papers. We give further information about this dichromatic formulation and its relationships with chromatic polynomials at the end of Section 3. See Figures 23, 24, 25.
@@
classical (and sometimes flat classical) crossings and virtual crossings. The classical crossings interact with one another via the Reidemeister moves and the virtual crossings interact via detour moves. In a detour move a consecutive sequence of virtual crossings of the same type can be excised from the diagram and the endpoints of this excision can be reconnected by any diagrammatic arc that intersects the rest of the diagram in a sequence of virtual crossings of this same type. In sections 4.1, 4.2, 4.3 and 4.4 we discuss this definition of multiple virtual knot theory and we give a sketch of the topological interpretation of this definition in the case of single virtual knot theory.
@@
Section 5 introduces generalizations of the quandle and the biquandle to multivirtual theory and discusses a number of examples, including a generalized Alexander polynomial.
@@ 2.3 Penrose Evaluation Formula
This formula recursive computation of the number of colorings for planar cubic graphs. As Penrose explains in his paper [49], the formula is a consequence of associating to each node in the graph an “epsilon” tensor
@@
(as we did before with a circle for virtual knot crossings but this is distinct from that method). The boxed flat crossing is a tensor depending on its four endpoints. It is zero unless the labels at the ends of a given straight segment in the crossed segments are the same. When these ends are the same we can label the crossed segments with two colors and as in Figure 10. Then the value of the tensor is if and if as shown in the Figure 10. The Penrose bracket is computed just as before and it satisfies the same formulas as before. The new crossings that occur in the expansion formula are standard crossings. Only the initial immersion crossings in the diagram are boxed. Examine the proof above and note that now every coloring will contribute to the state sum.
@@
The dotted crossing is regarded as a coloring vertex that demands that all four lines incident to this vertex have the same color. The ordinary flat crossing is taken as usual to indicate that each crossing line has a color independent of the other line. In the second equation we have replaced the ordinary flat crossing with a circled flat crossing. This means the same thing, and will be used later to discriminate these two types of virtual crossing as boxed and circled.
@@ 3 New Penrose Evaluations and Polynomials
a wiggly glyph are colored differently. We use the color set of colors described above.
@@
Figure 26 shows the simplest translation of perfect matching graphs to knot diagrams with virtual crossings. In this translation, we replace each perfect matching edge with a knot diagrammatic crossing so that a counterclockwise turn of the overcrossing arc sweeps the space previously occupied by the matching edge. With this convention we can articulate a coloring rule for the arcs of the knot or link diagram that corresponds directly to the rules for coloring the perfect matching. As the figure shows, the link diagram corresponding to a perfect matching graph may have virtual crossings if the original graph had such crossings, and the link diagram or knot diagram may be topologically knotted or unknotted depending on the distribution of the perfect matching edges. Virtual detour moves (See section 4 for the definition.) do not affect coloring of the knot or link diagram, but topological equivalence of knots (Reidemeister moves) can change the coloring properties. In the figure eight knot diagram at the bottom of the figure, we see that this diagram is the translation of a graph with a perfect matching that is not even. Consequently, all three colors appear on the arcs in the the three-coloring of the figure eight knot diagram. In the translation of the Petersen graph (with perfect matching) to a virtual link diagram, we see that the link consists in a small ring and a larger diagram with virtual crossings. Removal of the ring gives the larger diagram alone, which is uncolorable, but contains an isthmus (up to virtual detour moves on the surface of the two-sphere). The inclusion of the ring stops the isthmus and retains uncolorability. With this translation of perfect matching graphs to link diagrams, we see that coloring theorems can be studied in new terms. Classical link diagrams correspond to planar perfect matching graphs and so classical link diagrams with no isthmus are colorable with three colors (via the above rules) is an equivalent statement of the four color theorem.
@@ 4.3 Multiple Virtual Knot Theory
Detour Axiom for Multiple Virtual Crossings. Each virtual crossing type detours over all the other types of virtual crossings and over the classical crossings.
@@
classical Reidemeister moves. For a given instantiation of MVKT, a set of virtual crossing types is chosen with a method for indicating them. Note that in one aoplication of the detour move, only one virtual crossing type will be used.
Thus if one excises a sequence of virtual crossings of a given type, then then the new connection will create only virtual crossings of that type. If the excision does not remove any virtual crossings, then the reconnection can use
@@
Proof. The proof has the same structure as the proof for virtual knot theory in [28]. It is known from the work of Waldhausen that the fundamental group and periperal subgroup as a pair classify classical knots and links. Thus if and are classical links with peripheral subgroup pairs and then is classically ambient istopic to if and only if the corresponding group pairs are isomorphic. The group pair of a classical link or a virtual link diagram can be given by a Wirtinger presentation read from the diagram in conjunction with longitudes (one for each link component) that is obtained by walking along a component and writing the product of the Wirtinger generators (corrsponding to oriented arcs of the diagram) that are underpassed in the course of the walk. These longitudes are not affected by detour moves. From this it follows that if is obtained from by a sequence of Reidemeister moves and virtual detour moves (in the sense of MVKT), then is isomorphic as a pair with If and are classical diagrams, it follows that they are classically equivalent. This completes the proof of the Theorem.
@@ 4.5 Generalized Bracket Polynomial
Figure 33 where the local regions of a crossing are labeled and so that the regions are swept by a counterclockwise turn of the overcrossing arc, and a smoothing is of type if the two regions are joined by the smoothing. Each state in the virtual bracket expansion is a diagram with multiple virtual crossings. For the generalized bracket, we take to be the multiple flat virtual class (i.e. the equivalence class of the state up to virtual detour moves) of the state with the caveat that any free disjoint circle in (or in a diagram equivalent to ) is evaluated as Thus we have that
@@
Figure 38 illustrates our rule for the nodal crossing. This is in accord with the usage for the nodal crossing that was explained in Section 3 in relation to evaluating the
@@
corresponding flat crossing. The polynomial has the usual bracket expansion at a crossing coupled with our expansion for the box virtual crossings as we have used them for the graph theory earlier in the paper. In the formulas below we use the topological coefficients for the bracket with
@@
The context for the chromatic bracket is multi-virtuals with two crossing types: circle and box and the rule that the box virtual crossing is expanded into twice the nodal crossing minus the circle virtual. This is the identical formalism to our coloring polynomial from the Section 3. Hence one expands the bracket until there are only collecctions of virtual loops.
@@
nodes. We illustrate this calculation on the virtual flat link in Figure 40.
@@
We call this the topological specialization of the bracket. Thus the topological specialization for the chromatic bracket is an invariant of
double multiple virtual knots and links and is a well defined Laurent polynomial. For arbitrary multiple virtuals the topological specialization of the generalized bracket is invariant under the R2 and R3 moves and takes values in the ring generated by the flat virtual classes of the states of the diagram with coefficients in the Laurent polynomials in
@@
In Figure 40 we illustrate double virtual flat calculations for the simple link using two virtual crossings, and another example with three components.
In Figure 41 shows that the expansion for the box virtual crossing necessarily has the coefficients and if it is to satisfy the simplest flat Reidemeister two detour move.
@@
the detour moves and flat Reidemeister three moves (also detour moves) for the doubled virtual crossings. These figures constitute the crucial lemma that this method of evaluation leads to a regular isotopy invariant for doubled virtual links.
@@ 5 Quandles and Biquandles for Multiple Virtual Knots and Links
each of its classical and virtual crossings. We call the MultiVirtual Quandle associated to the link
Proposition. The Multivirtual Quandle associated with a multivirtual diagram is invariant under the classical Reidemeister moves and all detour moves.
Hence is an invariant of multivirtual links.
@@
The invariance under detours is verified in Figure 52. This completes the proof of the proposition.
In Figure 53 we illustrate the quandle coding of the double virtual trefoil , a multivirtual knot with one classical crossing, one round virtual crossing and one box virtual crossing. With the labelling as shown in the figure, and virtual free generators and we have that has generators and one relation
@@
To facilitate calculation Figure 54 shows the defining relations for the Generalized Alexander Quandle where the quandle is a module over the ring of integer-coefficient Laurent polynomials in the variable and in a set of commuting Laurent polynomial variables corresponding to the multiple virtual crossing types. In the figure we illustrate this quandle for just the two virtual types (round and box) with variables and respectively. Then we have
@@
multiple virtual knot Thus the knot is non-trivial and we have shown that its quandle, Alexander module and polynomial are all non-trivial.
@@
In Figure 56 we illustrate the quandle relations for two flat states that can appear in the expansion of knots and links by the generalized bracket. The first example is the state

### v2

@@ Abstract
We generalize virtual knot theory to multi-virtual knot theory where there are a multiplicity of virtual crossings. Each virtual crossing type can detour over the other virtual crossing types, and over classical or immersed crossings. New invariants of multi-virtual knots and links are introduced and new problems that arise are described. We show how the extensions of the Penrose coloring evaluation for trivalent plane graphs and our generalizations of this to non-planar graphs and arbitrary numbers of colors acts as a motivation for the construction of the multi-virtual theory.
@@ 1 Introduction
generalizations. Here we indicate a circled virtual crossing in the Penrose expansion. Sometimes the self-intersection is written without the circle, but we explain (two paragraphs below) why we have to indicate it. One can prove that is the number of proper three-colorings of a planar graph
Our approach to the Penrose bracket and its generalizations is based on the concept of a formation due to G. Spencer-Brown [53]. See Figure 17 for an illustration of a formation. In the initial case of three colors, we choose to illustrate directly the colors red (r) and blue (b) and to indicate the third color purple (p) as a superposition of red and blue. A formation is a planar drawing using red and blue Jordan curves so that the red curves are disjoint from one another, and the blue curves are disjoint from one another, but red and blue curves can share a finite set of disjoint intervals. The resulting drawing can be interpreted as a trivalent graph with a proper three coloring. The edges colored purple are in correspondence with the intervals in the drawing shared by red and. blue. An examination of the figure should clarify this statement. In the figure the corresponding trivalent graph is shown and the edges of the graph that correspond to shared red and blue are labeled with a slash. These purple edges in the coloring correspond to a perfect matching in the trivalent graph. This illustrates how a proper coloring gives rise to an even perfect matching (See Section 2). Conversely, if one has a properly colored trivalent graph, then there is a formation that corresponds to the coloring. The red and blue curves are regarded as crossing one another when one curve starts on one side of the other, shares an interval, and continues on the other side of the other curve.
@@
The dotted crossing is regarded as a coloring vertex that demands that all four lines incident to this vertex have the same color. When a boxed or circled crossing has colors assigned to its arcs it can be regarded as two straight arcs that cross one another at the circled or boxed intersection. Each straight arc has one color. The two colors may be the same or different. These interpretations are important for understanding the relationships with colorings, but are not used in the formal expansions of the formulas themselves, except at the very end. The straight arcs of the circled or boxed crossing can be colored with two colors that are different or the same. The technical reasons why this method adjusts the Penrose bracket count are reviewed in Section 2. Here it is important to point out that this second recursion is defined on the states of the Penrose bracket expansion. Each state consists in a collection of planar Jordan curves intersecting in round and square virtual crossings. The above evaluation is equivalent to a graphical contraction deletion algorithm for a graph where each state loop corresponds to a node of and two nodes are connected by an edge only when their loops touch at a boxed virtual crossing. Round virtual crossings do not affect the graph Then the above relation becomes the contraction deletion relation shown below.
@@
This means that we evaluate a dichromatic polynomial [37, 25, 17] for each state, and that the generalized Penrose evaluation is a sum with coefficients over these dichromatic evaluations. This is a useful direction of reformulation and will be followed up in subsequent papers. We give further information about this dichromatic formulation and its relationships with chromatic polynomials at the end of Section 3. See Figures 23, 24, 25.
@@
denote an edge of the graph associated with a boxed virtual crossing, as explained above.
@@
classical (and sometimes immersed) crossings and virtual crossings. The classical crossings interact with one another via the Reidemeister moves and the virtual crossings interact via detour moves. In a detour move a consecutive sequence of virtual crossings of the same type can be excised from the diagram and the endpoints of this excision can be reconnected by any diagrammatic arc that intersects the rest of the diagram in a sequence of virtual crossings of this same type. In sections 4.1, 4.2, 4.3 and 4.4 we discuss this definition of multiple virtual knot theory and we give a sketch of the topological interpretation of this definition in the case of single virtual knot theory.
@@
Section 5 introduces generalizations of the quandle and the biquandle to multi-virtual theory and discusses a number of examples, including a generalized Alexander polynomial.
@@ 2.3 Penrose Evaluation Formula
This formula gives a recursive computation of the number of colorings for planar cubic graphs. As Penrose explains in his paper [49], the formula is a consequence of associating to each node in the graph an “epsilon” tensor
@@
(as we did before with a circle for virtual knot crossings but this is distinct from that method). The boxed crossing is a tensor depending on its four endpoints. It is zero unless the labels at the ends of a given straight segment in the crossed segments are the same. When these ends are the same we can label the crossed segments with two colors and as in Figure 10. Then the value of the tensor is if and if as shown in the Figure 10. The Penrose bracket is computed just as before and it satisfies the same formulas as before. The new crossings that occur in the expansion formula are standard crossings. Only the initial immersion crossings in the diagram are boxed. Examine the proof above and note that now every coloring will contribute to the state sum.
@@
The dotted crossing is regarded as a coloring vertex that demands that all four lines incident to this vertex have the same color. The ordinary crossing is taken as usual to indicate that each crossing line has a color independent of the other line. In the second equation we have replaced the ordinary crossing with a circled crossing. This means the same thing, and will be used later to discriminate these two types of virtual crossing as boxed and circled.
@@ 3 New Penrose Evaluations and Polynomials
a wiggly glyph are colored differently. Thus the state evaluation is a chromatic polynomial of the graph associated with the state
by assigning a node to each loop in the state, and an edge to each pair of loops that share a wiggly line (indicating that the loops are colored differently). We use the color set of colors described above.
Remark.. The matching polynomial can be reformulated as a Penrose-type expansion in the form shown below, where it is given that loops that share a dark dotted intersection must be colored by the same color. Lines without wiggly lines or dotted intersections can be colored with the same color or with different colors. This formula is of interest, as it gives another solution to the problem to give a Penrose type formula that counts colors for all trivalent graphs. Consequences of this point of view will be studied in subsequent work. Note that for this formulation is any trivalent graph.
@@
Figure 26 shows the simplest translation of perfect matching graphs to knot diagrams with virtual crossings. In this translation, we replace each perfect matching edge with a knot diagrammatic crossing as illustrated at the top of the figure. In reference to the figure, note that the crossing edges and are paired and declared unequal, while the crossing edges and are paired and it is delcared that either and , or and This matches the corresponding condition for the perfect matching edge. With this convention we can articulate a coloring rule for semi-arcs of the knot or link diagram that corresponds directly to the rules for coloring the perfect matching. Here the semi-arcs of diagram are defined by the convention shown in Figure 26. Color can change at an over-crossing as indicated in that figure.
As the figure shows, the link diagram corresponding to a perfect matching graph may have virtual crossings if the original graph had such crossings, and the link diagram or knot diagram may be topologically knotted or unknotted depending on the distribution of the perfect matching edges. Virtual detour moves (See section 4 for the definition.) do not affect coloring of the knot or link diagram, but topological equivalence of knots (Reidemeister moves) can change the coloring properties. In the figure eight knot diagram at the bottom of the figure, we see that this diagram is the translation of a graph with a perfect matching that is not even. Consequently, all three colors appear on the arcs in the the three-coloring of the figure eight knot diagram. In the translation of the Petersen graph (with perfect matching) to a virtual link diagram, we see that the link consists in a small ring and a larger diagram with virtual crossings. Removal of the ring gives the larger diagram alone, which is uncolorable, but contains an isthmus (up to virtual detour moves on the surface of the two-sphere). The inclusion of the ring stops the isthmus and retains uncolorability. With this translation of perfect matching graphs to link diagrams, we see that coloring theorems can be studied in new terms. Classical link diagrams correspond to planar perfect matching graphs and so classical link diagrams with no isthmus are colorable with three colors (via the above rules) is an equivalent statement of the four color theorem.
@@ 4.3 Multiple Virtual Knot Theory
Detour Axiom for Multiple Virtual Crossings. Each virtual crossing type detours over all the types of virtual crossings including itself, and each virtual crossing type detours over the classical crossings.
@@
classical Reidemeister moves. Note that the virtual Reidemeister moves as shown in Figure 27 are consequences of the generalized detour move. The figure illustrates virtual Reidemeister moves for the case of one type of virtual and in the generalized theory there are such moves for each type of virtual crossing. For a given instantiation of MVKT, a set of virtual crossing types is chosen with a method for indicating them. Note that in one application of the detour move, only one virtual crossing type will be used.
Thus if one excises a sequence of virtual crossings of a given type, then the new connection will create only virtual crossings of that type. If the excision does not remove any virtual crossings, then the reconnection can use
@@
Proof. The proof has the same structure as the proof for virtual knot theory in [28]. It is known from the work of Waldhausen that the fundamental group and peripheral subgroup as a pair classify classical knots and links. Thus if and are classical links with peripheral subgroup pairs and then is classically ambient istopic to if and only if the corresponding group pairs are isomorphic. The group pair of a classical link or a virtual link diagram can be given by a Wirtinger presentation read from the diagram in conjunction with longitudes (one for each link component) that is obtained by walking along a component and writing the product of the Wirtinger generators (corrsponding to oriented arcs of the diagram) that are underpassed in the course of the walk. These longitudes are not affected by detour moves. From this it follows that if is obtained from by a sequence of Reidemeister moves and virtual detour moves (in the sense of MVKT), then is isomorphic as a pair with If and are classical diagrams, it follows that they are classically equivalent. This completes the proof of the Theorem.
@@ 4.5 Generalized Bracket Polynomial
Figure 33 where the local regions of a crossing are labeled and so that the regions are swept by a counterclockwise turn of the overcrossing arc, and a smoothing is of type if the two regions are joined by the smoothing. Each state in the virtual bracket expansion is a diagram with multiple virtual crossings. For the generalized bracket, we take to be the multiple virtual class (i.e. the equivalence class of the state up to virtual detour moves) of the state with the caveat that any free disjoint circle in (or in a diagram equivalent to ) is evaluated as Thus we have that
@@
Figure 38 illustrates our rule for the nodal crossing. This is in accord with the usage for the nodal crossing that was explained in Section 2.3 in relation to evaluating the
@@
corresponding crossing. The polynomial has the usual bracket expansion at a crossing coupled with our expansion for the box virtual crossings as we have used them for the graph theory earlier in the paper. In the formulas below we use the topological coefficients for the bracket with
@@
The context for the chromatic bracket is multi-virtuals with two crossing types: circle and box and the rule that the box virtual crossing is expanded into twice the nodal crossing minus the circle virtual. This is the identical formalism to our coloring polynomial from the Section 3. Hence one expands the bracket until there are only collections of virtual loops.
@@
nodes. We illustrate this calculation on the virtual diagram in Figure 40.
@@
We call this the topological specialization of the bracket. For arbitrary multiple virtuals the topological specialization of the generalized bracket is invariant under the R2 and R3 moves and takes values in the ring generated by the virtual classes of the states of the diagram with coefficients in the Laurent polynomials in
@@
In Figure 40 we illustrate double virtual calculations for the simple link using two virtual crossings, and another example with three components.
Figure 41 shows that the expansion for the box virtual crossing necessarily has the coefficients and if it is to satisfy the simplest virtual detour move.
@@
the detour moves and virtual Reidemeister three moves (also detour moves) for the doubled virtual crossings. These figures constitute the crucial lemma that this method of evaluation leads to a regular isotopy invariant for doubled virtual links.
@@ 5 Quandles and Biquandles for Multiple Virtual Knots and Links
Remark. Sometimes it is convenient to denote quandle operations by exponent notation. Thus we can write
and
It is understood that the “bar” in is strictly notational and does not designate a separate operation on This notation is used in Figure 60 and in the next section on biquandles.
@@
each of its classical and virtual crossings. We call the Multi-Virtual Quandle associated to the link
Proposition. The Multi-virtual Quandle associated with a multi-virtual diagram is invariant under the classical Reidemeister moves and all detour moves.
Hence is an invariant of multi-virtual links.
@@
The invariance under detours is verified in Figure 52, Figure 53 and Figure 54. Variations on the illustrated cases are left to the reader. This completes the proof of the proposition.
Remark. Note that one may wish to examine how these generalized quandles behave under the forbidden moves of Figure 29 and Figure 56. Note that in Figure 56 we illustrate the multiplicity of forbidden moves that arise in the presence of many virtual crossing types. When we say that moves are “forbidden” in a given theory, this does not mean that an invariant for that theory might not
have the same values or be invariant under these moves. The maxim that corresponds to axioms for moves is that “what is not allowed is forbidden”. We emphasize certain forbidden moves to remind ourselves and the reader that these moves are not consequences of the axioms and so invariants are not required to respect these moves. In the case of the forbidden moves under consideration, the bracket polynomials we have already discussed show that these moves are not consequences of the given axiomatic moves. It is often the case that quandles we define are also not invariant under these moves. In detail, see Figure 55 where we see that for a forbidden move of type to occur at the quandle level, the virtual automorphism representative would have to commute with an arbitrary element of the quandle. Since we only require the automorphisms for different virtual crossings to commute with one another, this extra relation will generically not occur.
Similarly, the requirement for the forbidden move of type is also an extra algebraic relation.
In Figure 57 we illustrate the quandle coding of the double virtual trefoil , a multi-virtual knot with one classical crossing, one round virtual crossing and one box virtual crossing. With the labelling as shown in the figure, and virtual free generators and we have that has generators and one relation
@@
To facilitate calculation Figure 58 shows the defining relations for the Generalized Alexander Quandle where the quandle is a module over the ring of integer-coefficient Laurent polynomials in the variable and in a set of commuting Laurent polynomial variables corresponding to the multiple virtual crossing types. In the figure we illustrate this quandle for just the two virtual types (round and box) with variables and respectively. Thus the ring is taken to be a ring of Laurent polynomials in three variables. The variables all commute with one another. Since elements of the quandle are modules over this ring, we can define quandle operations through linear multiplication by the elements of the ring. Then we have
@@
multiple virtual knot Thus the knot is non-trivial since we have shown that its polynomial is non-trivial. The non-triviality of the polynomial implies that the quandle from which it was computed is non-trivial and that the module that represents this quandle is non-trivial.
@@
In Figure 60 we illustrate the quandle relations for two virtual states that can appear in the expansion of knots and links by the generalized bracket. The first example is the state